OpenStudy (anonymous):

Which expression is equivalent to (√2 - 3√5)(√2 - √7) a. 2 - √14 - 3√10 + 3√35 b. 2 + √14 + 3√10 - 3√35 c. 2 - √7 + 3√10 - 3√35 d. 2 + √7 - 3√10 - 3√35 Thanks for those who answer! Please show the steps to your answer!

OpenStudy (anonymous):

I know -3√5*(-√7) = + 3√35 √2 * (- √17) = -√14 -3√5*√2 = -3√10 The only part I'm stuck on it how to get the

OpenStudy (anonymous):

2 in front of the equations

OpenStudy (shamim):

√2*√2=?